OpenStudy (anonymous):

In covering a distance of 30 km, Abhay takes 2 hours more than Sameer. If Abhay doubles his speed, then he would take 1 hour less than Sameer. Abhay's speed is: A. 5 kmph B. 6 kmph C. 6.25 kmph D. 7.5 kmph

OpenStudy (anonymous):

Hi \[\Delta x=30Km \rightarrow \Delta t=\Delta x/v \rightarrow \Delta t a=30/va , \Delta t s=30/vs \] \[30/va=2+30/vs , 30/2va=30/vs-1\rightarrow 30/2va=3 \rightarrow va=5kmph\]